# Efficient Exact Algorithms for Minimum Covering of Orthogonal Polygons with Squares

### v1

@@ 4.4 Polynomial Time Algorithm: Building up Partial Solutions
If the final set of rec-packs from Algorithm 4 is $\mathcal{R}$, then the running time is $\mathcal{O}(|\mathcal{R}|\cdot(n^{2}+|\mathcal{R}|n)^{4}))$.
Proof 4.12.
By Lemma 4.9, we know that each while loop iteration takes at most $\mathcal{O}((n^{2}+|\mathcal{R}|n)^{4})$ time. As the size of $\mathcal{R}$ increases at least by one at each iteration, the number of while loop iterations must be bounded by $|\mathcal{R}|$, the size of the final rec-pack minimum cover. Therefore, the total time for all while loop iterations is $\mathcal{O}(|\mathcal{R}|\cdot(n^{2}+|\mathcal{R}|n)^{4}))$. As there is only a total overhead of $\mathcal{O}(|\mathcal{R}|)$ outside the while loop, the total running time of Algorithm 4 becomes $\mathcal{O}(|\mathcal{R}|\cdot(n^{2}+|\mathcal{R}|n)^{4}))$.
Remark 4.13.
If we only extended our partial solutions by unambiguous squares (and not generate rec-packs from a seed), we would finally have a set of trivial rec-packs (squares) $\mathcal{S}$ which is a minimum cover of $P$ with OPCS $(P)=|\mathcal{S}|$. This would give us an output sensitive algorithm of $\mathcal{O}(|\mathcal{S}|\cdot(n^{2}+|\mathcal{S}|n)^{4}))$
We now proceed to bound the size of the final set $\mathcal{R}$ of rec-packs in Algorithm 4.
Lemma 4.14.
If $\mathcal{R}$ be the final set of rec-packs in Algorithm 4, then $|\mathcal{R}|=\mathcal{O}(n^{2})$
Proof 4.15.
Let $\mathcal{R}_{g}\subseteq\mathcal{R}$ be the set of rec-packs generated from a seed, and let $\mathcal{R}_{u}\subseteq\mathcal{R}$ be the set of unambiguous squares appearing as trivial rec-packs in $R$. Therefore, $\mathcal{R}=\mathcal{R}_{g}\cup\mathcal{R}_{u}$.
@@
For any set of squares $\mathcal{S}$ such that $\mathcal{E}(\mathcal{R}_{g})\cup\mathcal{S}$ covers $P$, we must have $|\mathcal{S}|\geq|\mathcal{R}_{u}|$.
Proof 4.16.
Since $\mathcal{R}$ is a minimum cover (Lemma 4.7), $\mathcal{E}(\mathcal{R})=\mathcal{E}(\mathcal{R}_{g})\cup\mathcal{R}_{u}$ is a minimum cardinality set of squares covering $P$. Therefore, for any set of squares $\mathcal{S}^{\prime}$ that covers $P$, we must have $|\mathcal{S}^{\prime}|\geq|\mathcal{E}(\mathcal{R}_{g})\cup\mathcal{R}_{u}|$. In particular, for any set of squares $\mathcal{S}$ such that $\mathcal{E}(\mathcal{R}_{g})\cup\mathcal{S}$ covers $P$, we must have $|\mathcal{E}(\mathcal{R}_{g})\cup\mathcal{S}|\geq|\mathcal{E}(\mathcal{R}_{g})\cup\mathcal{R}_{u}|$. This means that $|\mathcal{S}|\geq|\mathcal{R}_{u}|$ (as $\mathcal{R}_{u}$ and $\mathcal{E}(\mathcal{R}_{g})$ are disjoint).
We now construct such a set of squares $\mathcal{S}$ of size $\mathcal{O}(n^{2})$. Let us draw a vertical line and a horizontal line through each vertex $v$ of $P$. Let us call them vertex induced lines of $P$. As $P$ has $n$ vertices, there are at most $n$ vertex induced vertical lines and $n$ vertex induced horizontal lines.
These vertex induced lines form a grid like structure resulting in rectangular grid cells formed between two consecutive vertex induced vertical lines and two consecutive vertex induced horizontal lines. Let $abcd$ be a such a rectangular grid cell formed by consecutive vertex induced vertical & horizontal lines, that lies in the interior (possibly overlapping with some of the polygon edges of $P$), as shown in Figure 8. Let $a$ be the top-left corner, $b$ be the top-right corner, $c$ be the bottom-right corner and $d$ be the bottom left corner. Without loss of generality, let the rectangle $abcd$ have its vertical sides no longer than the horizontal sides (i.e. length is along the $x$-axis). We now present a method to cover the entire rectangular region $abcd$ (possibly covering some more portion of the interior of the polygon) with at most $5$ more valid squares (if we consider $\mathcal{R}_{g}$ to be already placed).
@@
Consider the two consecutive vertex induced vertical lines $l_{1},l_{2}$, one passing through $a,d$ and the other passing through $b,c$. There cannot be any vertex $v$ of $P$ lying strictly between these two lines (otherwise, these would not be adjacent vertex induced lines). As $abcd$ lies inside $P$, the following must hold (Figure 9(a)):
@@
  • There must be a horizontal polygon edge $e_{2}$ of $P$ that intersects both $l_{1}$ and $l_{2}$ (possibly at its endpoint), and that either overlaps with the line segment $cd$ or lies under $cd$.
@@
Let $\delta$ denote the vertical distance between $e_{1}$ and $e_{2}$. Further, if $\overline{e_{i}}$ is the length of an edge $e_{i}$, then $\overline{e_{1}},\overline{e_{2}}\geq\overline{ab}$ ($e_{1}$, $e_{2}$ intersect both $l_{1}$, $l_{2}$). We now discuss the following cases.
@@
  • Case I: $5\delta\geq\overline{ab}$. In this case, we can draw at most three $\delta\times\delta$ squares which cover the entire rectangular region enclosed by $e_{1}$, $e_{2}$, $l_{1}$, $l_{2}$ and hence covers $abcd$ (Figure 9(b)). Therefore $abcd$ can be covered by at most $3$ squares.
@@
  • Case II: $5\delta<\overline{ab}$. In this case, there must be a (unique) strip $Y$ that is enclosed within $e_{1}$, $e_{2}$ (Section 2). $Y$ must have an aspect ratio of more than $5$. Let $S_{0}\in\mathcal{R}$ be the first $\delta\times\delta$ square placed that covers some portion in $Y$ when Algorithm 4 is run on $P$ (there would always be a $\delta\times\delta$ square placed in $Y$ as no square with smaller dimension can generate a rec-pack as a seed that overlaps with $Y$, and all maximal squares covering any block which is at a distance of more than $\delta$ from the left & right boundary of $Y$, have dimension $\delta\times\delta$). Let $S_{l}$ be the left most $\delta\times\delta$ square lying inside $Y$ and $S_{L}$ be the square obtained by reflecting $S_{l}$ about its right side. Similarly, let $S_{r}$ be the right most $\delta\times\delta$ square lying inside $Y$ and $S_{R}$ be the square obtained by reflecting $S_{r}$ about its left side (Figure 10). Since the aspect ratio of $Y$ is more than $5$, $S_{l}$,$S_{L}$,$S_{R}$,$S_{r}$ are pairwise non-overlapping $\delta\times\delta$ squares that inside Y. Moreover since $Y$ is a strip with $\delta$ as the length of the shorter side, there cannot be a valid square that covers some region outside $Y$ as well as some region inside $Y$ which is not covered by $S_{l}$,$S_{r}$. Consider the portion $Y_{l}$ of $Y$ lying to the left hand side of $S_{0}$. Either $Y_{l}$ is entirely covered by $S_{l}$ and $S_{L}$, or the horizontal distance between the left side of $S_{0}$ and the left side of $Y_{l}$ is more than $2\delta$. Even if the second case occurs, then Algorithm 1 must generate $R_{l}\in\mathcal{R}_{g}$ as the rec-pack lying to the left of the seed $S_{0}$ which intersects with $S_{L}$ (as $S_{0}$ is the first square placed in the strip $Y$ and hence generates a rec-pack $R_{l}$ with the left side at most $2\delta$ distance from the left side of $Y$, and hence $R_{l}$ cannot intersect with any square or re-pack in the partial solution found before $S_{0}$ was added, Section 4.3). This means $Y_{l}$ is entirely covered by $S_{l}$, $S_{L}$ and possibly some rec-pack in $\mathcal{R}_{g}$. Similarly, the portion $Y_{r}$ of $Y$ to the right of $S$ is entirely covered by $S_{r}$, $S_{R}$ and possibly some rec-pack in $\mathcal{R}_{g}$. Therefore the entire region of $Y$ (and hence $abcd$) is covered by $\mathcal{R}_{g}$ and $5$ more squares, $\{S_{0},S_{l},S_{L},S_{R},S_{r}\}$.
@@
If we repeat the same process for each rectangular region $abcd$ formed by consecutive vertex induced lines, then we can cover up the entire polygon $P$ using just $5$ more squares for each such rectangular region lying inside $P$ (along with the rec-packs in $R_{g}$). However, the total number of such rectangular region is $\mathcal{O}(n^{2})$. Therefore, if $\mathcal{S}$ is the set of all such squares (at most $5$ of them per rectangular region $abcd$), then $|\mathcal{S}|=\mathcal{O}(n^{2})$ and $\mathcal{E}(\mathcal{R}_{g})\cup S$ is a set of squares covering the entire polygon $P$. As discussed in 1, we have $|\mathcal{R}_{u}|\leq|\mathcal{S}|=\mathcal{O}(n^{2})$.
@@
Moreover, as for each unambiguous square $S$, Algorithm 4 adds at most four rec-packs with seed $S$, the number of rec-packs generated from a seed can be at most $4$ times the number of unambiguous squares. Therefore $|\mathcal{R}_{g}|\leq 4|\mathcal{R}_{u}|=\mathcal{O}(n^{2})$.
Hence, $|\mathcal{R}|=|\mathcal{R}_{g}|+|\mathcal{R}_{u}|=\mathcal{O}(n^{2})$.
This allows us to derive the algorithmic running time purely in terms of $n$.
@@
For an orthogonal polygon $P$ with $n$ vertices, Algorithm 4 outputs $\textsc{OPCS}(P)$ in $\mathcal{O}(n^{14})$ time.
Proof 4.18.
The correctness of the algorithm is due to Lemma 4.7. For the analysis of the running time of Algorithm 4, we substitute the bound of $|\mathcal{R}|=\mathcal{O}(n^{2})$ obtained in Lemma 4.14 in the time complexity of Lemma 4.11. Therefore,
[…]
$=\mathcal{O}(n^{2}\cdot(n^{3})^{4})=\mathcal{O}(n^{14})$
This completes the proof.
@@
To report the solution instead of just the count, we could return the set $\mathcal{R}$ of rec-packs in Algorithm 4. Here, rec-packs are used to efficiently encode multiple squares (potentially exponentially many) into constant sized information.
Remark 4.20.
Here we assumed that arithmetic operations take constant time. However, even if each arithmetic operation took polynomial time, the total runtime would still be polynomial as there are only polynomially many arithmetic operations done.
@@ 5 Improved Algorithm for Orthogonal Polygons with $k$ Knobs
In this section, we design an algorithm for $p$-OPCS where along with the number $n$ of vertices of the input orthogonal polygon we also have the parameter $k$, which is the number of knobs (Definition 2.4) in $P$. This algorithm works better than Algorithm 4 when are input instances are such that $k=o(n^{13/14})$.
First, we define a special structure called separating squares and crucially use it to construct our recursive algorithm. We solve the base cases of the recursive algorithm using Algorithm 4.
@@ 5.1 More on Structure of Minimum Covering
We define our crucial structure of a separating maximal corner square (or simply, a separating square) and use the definition to further explore properties of non-knob convex vertices (Definition 2.5).
Definition 5.1 (Separating Maximal Corner Square).
Let $P$ be an orthogonal polygon. For a convex vertex $v$, $MCS(v)$ is said to be a separating maximal corner square or simply a separating square if the region inside $P$ but not inside $MCS(v)$ is not a connected region (Figure 11).
@@
Proof 5.3.
If $v_{i}$ is a non-knob convex vertex, then any curve lying in $P$ and having end points at $v_{i+2}$ and $v_{i-2}$ must intersect $MCS(v_{i})$ and hence $MCS(v_{i})$ must be a separating square separating $v_{i-2}$ and $v_{i+2}$.
@@
Given an orthogonal polygon $P$, we can find a non-knob convex vertex in $\mathcal{O}(n)$ time (if it exists) by a simple check on all vertices (or report no non-knob convex vertex exist).
We now use this to recursively obtain simpler instances from an input instance $P$.
@@ 5.2 Recursion with Separating Squares
Recall that a separating square separates an input orthogonal polygon $P$ into unconnected uncovered regions. We will construct two polygons from these uncovered polygons which still preserves the information about $\textsc{OPCS}(P)$. First, we need the following terminology.
@@
Given an orthogonal polygon $P$, let $S$ be a (maximal) separating square which is a maximal square due to a non-knob convex vertex of $P$. We classify the set of connected components of $P$ that are separated by $S$ as follows,
@@
Let $\mathcal{Q}^{\prime}=\{Q_{t},Q_{b},Q_{r},Q_{l},Q_{tr},Q_{br},Q_{tl},Q_{bl},Q_{trb},Q_{tlb},Q_{rbl},Q_{rtl}\}$ and let $\mathcal{Q}=\{Q\in\mathcal{Q}^{\prime}|Q\text{ is nonempty}\}$.
@@
Given an orthogonal polygon $P$, let $S$ be a (maximal) separating square which is a maximal square due to a non-knob convex vertex of $P$ and $\mathcal{Q}$ be as defined in 5.5. Then the following must hold true.
  • $\forall Q\in\mathcal{Q}$, $S\cup Q$ is a connected orthogonal polygon without holes.
  • $\text{{OPCS}}(P)=\big{(}\sum\limits_{Q\in\mathcal{Q}}\text{{OPCS}}(S\cup Q)\big{)}-(|\mathcal{Q}|-1)$
Figure 12: $\text{{OPCS}}(P)=\big{(}\sum\limits_{Q\in\mathcal{Q}}\text{{OPCS}}(S\cup Q)\big{)}-(|\mathcal{Q}|-1)$ with $\mathcal{Q}=\{Q_{l},Q_{rtl},Q_{r},Q_{bl},Q_{b}\}$ in this case
Proof 5.7.
@@
To see that for all $Q\in\mathcal{Q}$, $S\cup Q$ is connected, it is sufficient to observe that we can find a curve in the strict interior of $S\cup Q$ from any point $t_{1}$ in its interior to any point $t_{2}$ in its interior; either lying inside a single connected component of $Q$ (if $t_{1},t_{2}$ are in that component), or through $S$ (if $t_{1},t_{2}$ are in different components of $Q$). Further, $Q\cup S$ cannot have holes as $P$ does not have holes.
Moreover, we observe that there cannot be a valid square of $P$ that covers two different points from distinct $Q,Q^{\prime}\in\mathcal{Q}$ (this would contradict the maximality of $S$). Also, in a minimum covering with $S$ as one of the squares, all other valid squares must cover at least one point from the interior of exactly one $Q\in\mathcal{Q}$ (otherwise this square would be completely inside $S$, hence redundant). Therefore if $P$ is covered using a set $\mathcal{S}$ of $C$ squares including $S$, we can cover $Q\cup S$ using the squares in $\mathcal{S}$ that cover some part of $Q$ and the square $S$ itself. If we do this for all $Q\in\mathcal{Q}$ this uses $C+(|\mathcal{Q}|-1)$ squares as $S$ is in this cover of all $|\mathcal{Q}|$ instances $Q\cup S$, but appear only once in $\mathcal{S}$ (other squares in $\mathcal{S}$ appear exactly once in a cover of $Q\cup S$ for exactly one $Q\in\mathcal{Q}$). Since we start with a minimum cover of $P$ and find a valid cover of all $Q\cup S$ polygons with exactly $(|\mathcal{Q}|-1)$ more squares in total, we get,
$\text{{OPCS}}(P)\geq\big{(}\sum\limits_{Q\in\mathcal{Q}}\text{{OPCS}}(S\cup Q)\big{)}-(|\mathcal{Q}|-1)$
Given any minimum covering of all $Q\cup S$, all such instances must contain $S$ (as $S$ is a maximum corner square in both). So we superimpose these coverings and delete all but one $S$ to get a covering of $P$. This time we started from a minimal covering of the individual $Q\cup S$ polygons to get a valid covering of $P$. Including this with the rest of the result, we obtain
$\text{{OPCS}}(P)=\big{(}\sum\limits_{Q\in\mathcal{Q}}\text{{OPCS}}(S\cup Q)\big{)}-(|\mathcal{Q}|-1)$
Remark 5.8.
$|\mathcal{Q}|\leq|\mathcal{Q}^{\prime}|=12$ in Lemma 5.6. Moreover, there are combinations (like $Q_{tr}$ and $Q_{trb}$) that cannot both be non-empty in $P$.
We now prove a crucial result that such a recursive step does not increase the number of knobs in each individual instance $Q\cup S$.
@@
Let $S$ be a (maximal) separating square which is a maximal square due to a non-knob convex vertex of an orthogonal polygon $P$ with $n$ vertices and $k$ knobs. Let $\mathcal{Q}$ be defined as in Definition 5.5. Then, $\forall Q\in\mathcal{Q},(Q\cup S)$ is an orthogonal polygon without holes with at most $n$ vertices and at most $k$ knobs. Moreover, any vertex in $Q\cup S$ which is a corner of $S$ is part of a knob in $Q\cup S$ along a side of $S$.
Proof 5.10.
Lemma 5.6 already proves that $\forall Q\in\mathcal{Q},(Q\cup S)$ is an orthogonal polygon without holes. Further, the number of vertices can only decrease because the only time a new vertex (vertex not in $P$) would be introduced is when $S$ already distributes the existing vertices of $P$ in each of $(Q\cup S),Q\in\mathcal{Q}$ (causing no total increase in the number of vertices in $(Q\cup S)$ than in $P$). We now prove that $\forall Q\in\mathcal{Q},Q\cup P$ has at most $k$ knobs.
Consider $Q\cup S$ for some $Q\in\mathcal{Q}$. Without loss of generality assume $Q$ is either $Q_{l}$ or $Q_{tl}$ or $Q{tlb}$ (all other cases have symmetric arguments). We now show that for all knobs in $Q\cup S$, there is a distinct knob in $P$ (which would show that $Q\cup S$ has at most $k$ knobs if $P$ has at most $k$ knobs. Clearly any (distinct) knob $(u1,u2)$ of $Q\cup S$ such that $u1$, $u_{2}$ are not corners in $S$, $(u_{1},u_{2})$ must be a (distinct) knob in $P$ (knob in $Q$, in particular). Hence we only need to consider knobs which $(u_{1},u_{2})$ in $Q\cup S$ such that either $u_{1}$ or $u_{2}$ is a vertex of $S$. Without loss of generality, we assume that $u_{1}$ is a corner in $S$.

### v2

@@ 5.4 Polynomial-time algorithm: building up partial solutions
If $\mathscr{R}$ be the final set of rec-packs in Algorithm 3, then $|\mathscr{R}|=\mathcal{O}(n^{2})$
Proof 5.12.
Let $\mathscr{R}_{g}\subseteq\mathscr{R}$ be the set of rec-packs generated from a seed, and let $\mathscr{R}_{u}\subseteq\mathscr{R}$ be the set of unambiguous squares appearing as trivial rec-packs in $\mathscr{R}$. Therefore, $\mathscr{R}=\mathscr{R}_{g}\cup\mathscr{R}_{u}$.
@@
For any set of squares $\mathscr{S}$ such that $\mathsf{ext}(\mathscr{R}_{g})\cup\mathscr{S}$ covers $P$, we must have $|\mathscr{S}|\geq|\mathscr{R}_{u}|$.
Proof 5.13 (Proof of 4).
Since $\mathscr{R}$ is a minimum cover (Lemma 5.7), $\mathsf{ext}(\mathscr{R})=\mathsf{ext}(\mathscr{R}_{g})\cup\mathscr{R}_{u}$ is a minimum cardinality set of squares covering $P$. Therefore, for any set of squares $\mathscr{S}^{\prime}$ that covers $P$, we must have $|\mathscr{S}^{\prime}|\geq|\mathsf{ext}(\mathscr{R}_{g})\cup\mathscr{R}_{u}|$. In particular, for any set of squares $\mathscr{S}$ such that $\mathsf{ext}(\mathscr{R}_{g})\cup\mathscr{S}$ covers $P$, we must have $|\mathsf{ext}(\mathscr{R}_{g})\cup\mathscr{S}|\geq|\mathsf{ext}(\mathscr{R}_{g})\cup\mathscr{R}_{u}|$. This means that $|\mathscr{S}|\geq|\mathscr{R}_{u}|$, as $\mathscr{R}_{u}$ and $\mathsf{ext}(\mathscr{R}_{g})$ are disjoint.
We now construct such a set of squares $\mathscr{S}$ of size $\mathcal{O}(n^{2})$, with $\mathsf{ext}(R_{g})\cup\mathscr{S}$ entirely covering $P$. Let us draw a vertical line and a horizontal line through each vertex $v$ of $P$. We will refer to these as vertex-induced lines of $P$. As $P$ has $n$ vertices, there are at most $n$ vertex-induced vertical lines and $n$ vertex-induced horizontal lines.
These vertex-induced lines form a grid-like structure consisting of rectangular grid cells. Each rectangular grid cell is formed between two consecutive vertex-induced vertical lines and two consecutive vertex-induced horizontal lines. Let $abcd$ be a such a rectangular grid cell, that does not lie outside $P$ (Figure 10). Let $a$ be the top-left corner, $b$ be the top-right corner, $c$ be the bottom-right corner and $d$ be the bottom-left corner. Without loss of generality, assume that the rectangle $abcd$ has its vertical sides are shorter than or the same as its horizontal sides.
Consider the set of rec-packs in $R_{g}$ to already be placed. We now present a method to cover the entire rectangular region $abcd$ (possibly covering some more portion of the interior of the polygon) with at most $5$ more valid squares.
@@
Consider the two consecutive vertex-induced vertical lines $l_{1},l_{2}$, passing through $a,d$ and passing through $b,c$ respectively. Notice that there cannot be any vertex $v$ of $P$ lying strictly between these two lines; otherwise the vertical vertex-induced line due to this vertex will make $l_{1}$ and $l_{2}$ not consecutive. As $abcd$ lies inside $P$, the following must hold. (Figure 11(a)):
@@
  • There must be a horizontal polygon edge $e_{2}$ of $P$ that intersects both $l_{1}$ and $l_{2}$ (possibly at its endpoint), and that either overlaps with the line segment $cd$ or lies below $cd$.
@@
If $\overline{e_{i}}$ denotes the length of the edge $e_{i}$, then $\overline{e_{1}},\overline{e_{2}}\geq\overline{ab}$, as $e_{1}$, $e_{2}$ intersect both $l_{1}$, $l_{2}$. Let $\delta$ denote the (vertical) distance between $e_{1}$ and $e_{2}$. We now discuss the following cases.
@@
  • Case I: $5\delta\geq\overline{ab}$. In this case, we can draw at most five $\delta\times\delta$ squares which cover the entire rectangular region enclosed by $e_{1}$, $e_{2}$, $l_{1}$, $l_{2}$ and hence covers $abcd$ (Figure 11(b)). Therefore $abcd$ can be covered by at most $5$ squares.
@@
  • Case II: $5\delta<\overline{ab}$. In this case, there must be a unique strip $Y$ that is enclosed within $e_{1}$, $e_{2}$ (Section 2). Moreover, $Y$ must have an aspect ratio of more than $5$. Let $S_{l}$ be the leftmost $\delta\times\delta$ square lying inside $Y$ and $S_{L}$ be the square obtained by reflecting $S_{l}$ about its right side. Similarly, let $S_{r}$ be the rightmost $\delta\times\delta$ square lying inside $Y$ and $S_{R}$ be the square obtained by reflecting $S_{r}$ about its left side (Figure 12). Since the aspect ratio of $Y$ is more than $5$, the squares $S_{l}$,$S_{L}$,$S_{R}$,$S_{r}$ are pairwise non-overlapping $\delta\times\delta$ squares lying inside $Y$. Let $Y^{\prime}$ denote the region in $Y$ not covered by $S_{l}$, $S_{L}$, $S_{R}$, $S_{r}$. Since $Y$ is a strip with $\delta$ as the length of the shorter side, any valid square that covers some region outside $Y$ and some region inside $Y$, must overlap with $S_{l}$ or $S_{r}$, and can never overlap with $S_{L}$ and $S_{R}$.
    Let $S_{0}\in\mathscr{R}$ be the first $\delta\times\delta$ square placed that covers some portion in $Y$ when Algorithm 3 is run on $P$. We look into the rec-packs generated when $S_{0}$ is considered as a seed (Algorithm 2); rec-packs generated are only to the right or to the left. Recall that we look at the longest rec-packs, and any square covering points outside strip $Y$ does not overlap with $S_{L}$, $S_{R}$. Therefore, $S_{0}$ along with the rec-packs generated in $\mathscr{R}_{g}$ with $S_{0}$ as a seed together cover $Y^{\prime}$, and possibly some portions of $S_{l}$, $S_{L}$, $S_{R}$, $S_{r}$. Hence, with all rec-packs in $R_{g}$ already placed, we only need $5$ more squares, $S_{0}$, $S_{l}$, $S_{L}$, $S_{R}$, $S_{r}$ to cover $Y$ (and therefore cover $abcd$).
@@
If we repeat the same process for each rectangular grid-cell $abcd$ formed by consecutive vertex-induced lines, then we can cover up the entire polygon $P$ using just $5$ more squares for each such rectangular region lying inside $P$ (along with the rec-packs in $R_{g}$). However, the total number of such rectangular grid-cells is $\mathcal{O}(n^{2})$. Therefore, if $\mathscr{S}$ is the set of all such squares (at most $5$ of them per rectangular region $abcd$), then $|\mathscr{S}|=\mathcal{O}(n^{2})$ and $\mathsf{ext}(\mathscr{R}_{g})\cup S$ is a set of squares covering the entire polygon $P$. As discussed in 4, we have $|\mathscr{R}_{u}|\leq|\mathscr{S}|=\mathcal{O}(n^{2})$.
@@
Moreover, as for each unambiguous square $S$, Algorithm 3 adds at most four rec-packs with seed $S$, the number of rec-packs generated from a seed can be at most $4$ times the number of unambiguous squares. Therefore $|\mathscr{R}_{g}|\leq 4|\mathscr{R}_{u}|=\mathcal{O}(n^{2})$.
Hence, $|\mathscr{R}|=|\mathscr{R}_{g}|+|\mathscr{R}_{u}|=\mathcal{O}(n^{2})$.
Putting together Lemma 5.9 and Lemma 5.11, we get that the total time complexity of Algorithm 3 is $\mathcal{O}(n^{2}\cdot n^{2}\cdot(n+n^{2})^{3})=\mathcal{O}(n^{10})$
@@
For an orthogonal polygon $P$ with $n$ vertices, Algorithm 3 solves OPCS in $\mathcal{O}(n^{10})$ time.
@@
To report the solution instead of just the count, we could return the set $\mathscr{R}$ of rec-packs in Algorithm 3. Here, rec-packs are used to efficiently encode multiple squares (potentially exponentially many) into constant sized information.
@@ 6 Improved Algorithm for Orthogonal Polygons with $k$ Knobs
In this section, we consider the $p$-OPCS problem, where our input polygon $P$, has $n$ vertices, and at most $k$ knobs (Definition 2.5), where $k$ is a parameter. We design an algorithm for $p$-OPCS that is more efficient than Algorithm 3, whenever the input instances are such that $k=o(n^{9/10})$.
First, we define a special structure called separating squares, and prove some of its structural results. We will crucially use this to construct our recursive algorithm.
@@ 6.1 More on structure of minimum covering
We define a separating square, and use the definition to further explore properties of non-knob convex vertices (Definition 2.6).
Definition 6.1 (Separating Square).
For a convex vertex $v$ of an orthogonal polygon $P$, $MCS(v)$ is said to be a separating square if the region inside $P$ but outside $MCS(v)$ is not a connected region (Figure 13).
@@
Proof 6.3.
If $v_{i}$ is a non-knob convex vertex, then any curve lying inside $P$, and having end points at $v_{i+2}$ and $v_{i-2}$ must intersect $MCS(v_{i})$ and hence $MCS(v_{i})$ must be a separating square separating $v_{i-2}$ and $v_{i+2}$.
@@
Given an orthogonal polygon $P$, we can find a non-knob convex vertex in $\mathcal{O}(n)$ time (or report that it does not exist) by a simple check on all vertices.
We now use this to recursively obtain simpler.
@@ 6.2 Recursion with separating squares
A separating square separates an input orthogonal polygon $P$ into unconnected uncovered regions. We will construct two or more polygons from these uncovered polygons which still preserves the information about $\textsc{OPCS}(P)$.
First, we define the following.
@@
Given an orthogonal polygon $P$, let $S$ be a maximal separating square which is a maximal square due to a non-knob convex vertex of $P$. We classify the set of connected components of $P$ that are separated by $S$ as follows,
@@
[…]
Please refer to Figure 14 for illustrations of some of these components.
@@
Given an orthogonal polygon $P$, let $S$ be a maximal separating square which is a maximal square due to a non-knob convex vertex of $P$. Let $\mathscr{Q}$ be as defined in Definition 6.5. Then the following must hold true.
  • $\forall Q\in\mathscr{Q}$, $S\cup Q$ is a connected orthogonal polygon without holes.
  • $\text{{OPCS}}(P)=\left(\sum\limits_{Q\in\mathscr{Q}}\text{{OPCS}}(S\cup Q)\right)-(|\mathscr{Q}|-1)$
Figure 14: $\text{{OPCS}}(P)=\left(\sum\limits_{Q\in\mathscr{Q}}\text{{OPCS}}(S\cup Q)\right)-(|\mathscr{Q}|-1)$ with $\mathscr{Q}=\{Q_{l},Q_{rtl},Q_{r},Q_{bl},Q_{b}\}$ in this case
Proof 6.7.
@@
To see that for all $Q\in\mathscr{Q}$, $S\cup Q$ is connected, it is sufficient to observe that we can find a curve in the strict interior of $S\cup Q$ from any point $t_{1}$ in its interior to any point $t_{2}$ in its interior; either lying inside a single connected component of $Q$ (if $t_{1},t_{2}$ are in that component), or through $S$ (if $t_{1},t_{2}$ are in different components of $Q$). Further, $Q\cup S$ cannot have holes as $P$ does not have holes.
Since $S$ is maximal, observe that there cannot be a valid square of $P$, that covers two different points from distinct $Q,Q^{\prime}\in\mathscr{Q}$. Also, in a minimum covering with $S$ as one of the squares, all other valid squares must cover at least one point from the interior of exactly one $Q\in\mathscr{Q}$ (otherwise this square would be completely inside $S$, hence redundant). Therefore if $P$ is covered using a set $\mathscr{S}$ of $C$ squares, where $S\in\mathscr{S}$, we can cover $Q\cup S$ using the squares in $\mathscr{S}$ that cover some part of $Q$ and the square $S$ itself. If we do this for all $Q\in\mathscr{Q}$ this uses $C+(|\mathscr{Q}|-1)$ squares, as $S$ is in this cover of all $|\mathscr{Q}|$ instances $Q\cup S$. But $S$ appears only once in $\mathscr{S}$. Since we start with a minimum cover of $P$ and find a valid cover of all $Q\cup S$ polygons with exactly $(|\mathscr{Q}|-1)$ more squares in total, we get,
$\text{{OPCS}}(P)\geq\left(\sum\limits_{Q\in\mathscr{Q}}\text{{OPCS}}(S\cup Q)\right)-(|\mathscr{Q}|-1)$
Now, given any minimum covering of all $Q\cup S$, all such instances must contain $S$ (as $S$ is a maximum corner square). So we superimpose these coverings and delete all but one copy of $S$ to get a covering of $P$. This time we started from a minimal covering of the individual $Q\cup S$ polygons to get a valid covering of $P$. Including this with the rest of the result, we obtain
$\text{{OPCS}}(P)=\left(\sum\limits_{Q\in\mathscr{Q}}\text{{OPCS}}(S\cup Q)\right)-(|\mathscr{Q}|-1)$
We now prove a crucial result: such a recursive step does not increase the number of knobs in each individual instance $Q\cup S$.
@@
Let $S$ be a maximal separating square which is a maximal square due to a non-knob convex vertex of an orthogonal polygon $P$ with $n$ vertices and $k$ knobs. Let $\mathscr{Q}$ be defined as in Definition 6.5. Then, for every $Q\in\mathscr{Q}$, $(Q\cup S)$ is an orthogonal polygon without holes with at most $n$ vertices and at most $k$ knobs. Moreover, any vertex in $Q\cup S$ which is a corner of $S$ is part of a knob in $Q\cup S$ that coincides with a side of $S$.
Proof 6.9.
Lemma 6.6 already proves that $\forall Q\in\mathscr{Q},(Q\cup S)$ is an orthogonal polygon without holes. Further, the number of vertices can only decrease because the only time a new vertex (vertex not in $P$) would be introduced is when $S$ already distributes the existing vertices of $P$ in each of $(Q\cup S),Q\in\mathscr{Q}$ (causing no total increase in the number of vertices in $(Q\cup S)$ than in $P$). We now prove that $\forall Q\in\mathscr{Q},Q\cup P$ has at most $k$ knobs.
Consider $Q\cup S$ for some $Q\in\mathscr{Q}$. Without loss of generality assume $Q$ is either $Q_{l}$ or $Q_{tl}$ or $Q{tlb}$ (all other cases have symmetric arguments). We now show that for all knobs in $Q\cup S$, there is a distinct knob in $P$ (which would show that $Q\cup S$ has at most $k$ knobs if $P$ has at most $k$ knobs. Clearly any (distinct) knob $(u1,u2)$ of $Q\cup S$ such that $u1$, $u_{2}$ are not corners in $S$, $(u_{1},u_{2})$ must be a (distinct) knob in $P$ (knob in $Q$, in particular). Hence we only need to consider knobs which $(u_{1},u_{2})$ in $Q\cup S$ such that either $u_{1}$ or $u_{2}$ is a vertex of $S$. Without loss of generality, we assume that $u_{1}$ is a corner in $S$.